2.49 You throw a small rock straight up from the edge of a highway bridge that crosses a river. The rock passes you on its way down, 6.00 s after it was thrown. What is the speed of the rock just before it reaches the water 28.0 m below the point where the rock left your hand? Ignore air resistance.
Answer #1

answer) first we have time of flight

T=2v/g

6s=2*v/9.8m/s2

v=9.8*6/2=29.4m/s

this is the speed when the rock passes you

now we have the formula

v2=u2+2gs

v2=29.42+2*9.8m/s2*28m

v=\sqrt{}1413.16=37.6m/s

so the answer is 37.6m/s or 38m/s.